# The Quantum Esscher Transform

### v1

@@ Abstract
1 Introduction
In probability and statistics, it is often important to find low relative-entropy distributions from a given fixed distribution. In addition, further constraints, the form and interpretation of which depend on the problem at hand, are frequently imposed on the target distribution.
An interesting example is the following: consider the process of inferring probability distributions from a set of measurement data. These data play the role of the constraints—they put restrictions on what the true distribution could be—and the available data may not suffice to uniquely determine a probability distribution. In this situation, a common approach is to invoke Jaynes’ maximum entropy principle (MaxEnt) [Jay57]. In essence, MaxEnt advocates that the selected distribution be the one that simultaneously maximizes entropy and satisfies the given constraints.
However, the situation becomes more nuanced if we already possess some knowledge of the system, say, a prior distribution. In such cases, a more refined strategy emerges: the minimum relative entropy principle. As expounded in [SJ80, OP07, ZTF13], this principle, regarded as a generalization of MaxEnt, operates by minimizing the distinguishability (characterized by the relative entropy) between the prior distribution and the distribution to be selected, while respecting the imposed constraints. This systematic approach to incorporating new data makes it fundamental in Bayesian statistics. The updating procedure results in the posterior distribution which reflects the most current understanding of the system in light of the observed data.
In the case when the measurement data is presented in the form of expectation values of selected random variables, the solution to the corresponding relative entropy minimization problem takes the form known as an Esscher Transform. Named after Swedish mathematician and economist Fredrik Esscher, who introduced the concept in 1932 in his work on risk theory [Esc32], the Esscher Transform, also known as ‘exponential tilting’ in statistics, and its various extensions have since then found many applications beyond minimizing relative entropy. Notable examples include option pricing (in mathematical finance) [${}^{+}$93], importance sampling (for rare-event simulation) [Sie76] and Lévy processes (in financial economics) [HS06]. More recently, it has also made inroads into machine learning [BSS23], in the context of empirical risk minimization.
@@ 1 Introduction
$\text{minimize}_{\sigma\geq 0}$ $S(\sigma\|\rho)$ (1.1)
@@
where $\rho$ is the a priori state and $H_{i}$, $i\in[d]$ are observables. Refer to Definition 2.4 for the precise formulation. In the first part of this work, we show the formal solution to this constrained optimization problem. The solution methodology is modelled after its classical predecessor, albeit with added technical intricacies to manage. The form of the corresponding solution then motivates us to define the quantum Esscher Transform, see Definition 2.8. The proof of the solution to the optimization problem is found in Theorem 2.5. The quantum Esscher Transform can be viewed as a generalization of the (classical) Esscher Transform, and indeed subsumes the latter as a special case. In the second part of this work, with an eye toward potential applications, we discuss the implementation of the quantum Esscher Transform on fault-tolerant quantum computers. Our algorithm is based on the modern techniques of block-encoding and the quantum singular value transformation (QSVT) [GSLW19, MRTC21]. As an input model we consider purifications of the density operator $\rho$ and block-encodings of the operators $H_{i}$. The main algorithm is Algorithm 1, whose complexity is discussed in Theorem 4.3. The quantum Esscher transform could find applications in quantum analogues of problems in statistics, machine learning, and finance.
1.1 Preliminaries and notation
@@
2 Quantum Esscher Transform
2.1 Esscher Transform
@@
The Esscher Transform is a map from and onto the space of probability mass/density functions, as $\mathcal{E}(f;\theta)=f_{\theta}$. In this work, we never invoke $\mathcal{E}$ and simply call $f_{\theta}$ the Esscher Transform of $f$, in the same spirit as the Fourier Transform. In the context of probability theory, let $(\Omega,\Sigma,P)$ be a probability space and $X:\Omega\longrightarrow\mathbb{R}^{d}$ a random $d$ dimensional vector. This setting motivates the equivalent definition (see Remark 2.3 below) of Esscher Transforms for measures/distributions.
Definition 2.1 (Esscher Transform for probability distributions).
@@
$\text{minimize}_{Q\in[0,1]^{|\Omega|}}$ $D(Q\|P)$ (2.1)
@@
Note that there are $d+1$ constraints on $Q$, hence in feasible, non-redundant cases we have $d+1\leq|\Omega|$. We have the following solution to the optimization problem.
Theorem 2.2.
Given a random vector $X:\Omega\longrightarrow\mathbb{R}^{d}$ and $m\in\mathbb{R}^{d}$ where $\min_{\omega\in\Omega}X_{i}(\omega)<m_{i}<\max_{\omega\in\Omega}X_{i}(\omega)$ for $i\in[d]$. There exists a unique solution $Q^{\star}$ to problem 2.1, given by
@@
where $\lambda^{\star}:=\operatorname*{argmin}_{\lambda\in\mathbb{R}^{d}}\mathbb{E}_{P}[e^{\lambda\cdot(X-m)}]$. Thus $Q^{\star}$ is the $(\lambda^{\star},X)$-Esscher Transform of $P$, see Definition 2.1.
The proof is elaborated in Appendix A.
Remark 2.3.
Let us comment on a subtlety. Above, we have called $Q^{\star}$ the Esscher Transform of $P$. Recall that the Esscher Transform as originally defined by Esscher pertains to probability mass/density functions instead of measures. Here we show that using the same terminology for probability measures is well-justified (at least for the case when $\Omega$ is discrete). The random variable $X$ induces from the probability measure $P$ the probability mass function $P_{X}(x):=P(X^{-1}(x))$ on $E:=X(\Omega)$. Assume we have, for probability measures $Q,P$ and random variable $X$, that
Then for the probability mass functions $Q_{X}$ and $P_{X}$ we have
$Q_{X}(x)=Q(X^{-1}(x))$ $=\sum_{\omega:X(\omega)=x}Q(\omega)$
$=\frac{e^{\theta\cdot x}P_{X}(x)}{\sum_{x\in E}e^{\theta\cdot x}P_{X}(x)},$
i.e., $Q_{X}$ is the Esscher Transform of $P_{X}$ as defined above.
2.2 Quantum version
2.2.1 Problem statement
Many entities in classical probability theory have meaningful generalizations in quantum theory. For example, sample spaces, probability distributions and random variables find their respective counterparts in Hilbert spaces, density operators and observables (the latter also include the former as special instances). The quantum counterpart of the relative entropy is the quantum relative entropy,
@@ 2.2.1 Problem statement
defined for density operators $\sigma,\rho$. As in the classical case, the definition of $S(\sigma\|\rho)$ imposes constraints on $\sigma$ and $\rho$ in order to have $S(\sigma\|\rho)<\infty$. Namely, $\operatorname{supp}(\sigma)\subseteq\operatorname{supp}(\rho)$ (see Chapter 11, [Wil13]) or equivalently, $\ker(\rho)\subseteq\ker(\sigma)$. Using terminology from measure theory, if this condition is satisfied we say $\sigma$ is absolutely continuous with respect to $\rho$ ($\sigma\ll\rho$). This is analogous to the absolute continuity between probability distributions in classical probability theory. Now we formally state the quantized version of Problem 2.1.
Problem 2.4.
Let $\mathcal{H}_{N}$ be an $N$-dimensional Hilbert space and $\rho\in\mathcal{D}(\mathcal{H}_{N})$ be a density operator. With $d\in\mathbb{N}$, for $i\in[d]$, let $H_{i}$ be an observable with $h_{i,\min}$ and $h_{i,\max}$ denoting its smallest and largest eigenvalue respectively. For $m\in\mathbb{R}^{d}$ with $h_{i,\min}<m_{i}<h_{i,\max}$, solve
$\text{minimize}_{\sigma\geq 0}$ $S(\sigma\|\rho)$ (2.2)
@@
2.2.2 Solution
Before delving into the solution, let us briefly comment on a few possible concerns. First, $S(\sigma\|\rho)$ requires taking the logarithm of $\rho$, which poses a problem if $\rho$ is not strictly positive definite. This issue is circumvented if, as mentioned above, $\ker(\rho)\subseteq\ker(\sigma)$. The analysis becomes relatively straightforward if we partition the Hilbert space $\mathcal{H}$ into suitable subspaces and examine $\sigma$ over them separately. To this end, we introduce the following notation. Let $\mathcal{G}$ be a subspace of $\mathcal{H}$. For $A\in\mathcal{L}(\mathcal{H})$, denote $A_{\mathcal{G}}:=\Pi_{\mathcal{G}}A\Pi_{\mathcal{G}}\in\mathcal{L}(\mathcal{G})$, where $\Pi_{\mathcal{G}}$ is the projector onto $\mathcal{G}$.
@@ 2.2.2 Solution
We have the following result, which partially resolves Problem 2.4:
Theorem 2.5.
The solution to Problem 2.4 takes the form
$\sigma^{\star}=\sigma_{\operatorname{supp}\rho}^{\star}\oplus\sigma_{\ker\rho}^{\star},$ (2.3)
@@
$\sigma_{\operatorname{supp}\rho}^{\star}=\frac{e^{\lambda^{\star}\cdot H_{\operatorname{supp}\rho}+\log\rho_{\operatorname{supp}\rho}}}{\operatorname{Tr}(e^{\lambda^{\star}\cdot H_{\operatorname{supp}\rho}+\log\rho_{\operatorname{supp}\rho}})}\qquad\text{and}\qquad\sigma_{\ker\rho}^{\star}=\mathbf{0}.$ (2.4)
@@
$\operatorname{Tr}\left(e^{\lambda^{\star}\cdot(H_{\operatorname{supp}\rho}-m)+\log\rho_{\operatorname{supp}\rho}}(H_{i,\operatorname{supp}\rho}-m_{i})\right)=0\;,i\in[d].$ (2.5)
Proof.
Step 1. First, for any candidate solution $\sigma$ we enforce $\ker\rho\subseteq\ker\sigma$. By Lemma 2.6, this implies $\sigma_{\ker\rho}=\mathbf{0}$ and furthermore enables the decomposition of $\sigma$ into a direct sum: $\sigma=\sigma_{\operatorname{supp}\rho}\oplus\sigma_{\ker\rho}$. With this decomposition, we can consider the trace of the operators over just the subspace $\operatorname{supp}\rho$. More specifically, $\operatorname{Tr}(\sigma H_{i})=\operatorname{Tr}(\sigma(\Pi_{\operatorname{supp}\rho}+\Pi_{\ker\rho})H_{i}(\Pi_{\operatorname{supp}\rho}+\Pi_{\ker\rho}))=\operatorname{Tr}(\sigma_{\operatorname{supp}\rho}H_{i,\operatorname{supp}\rho})$ [^1] Recall that for any $A\in\mathcal{L}(\mathcal{H})$, $\ker A\oplus\operatorname{supp}A=\mathcal{H}$, so $\Pi_{\ker A}+\Pi_{\operatorname{supp}A}=I$. and
$S(\sigma\|\rho)$ $=\operatorname{Tr}\{\sigma_{\operatorname{supp}\rho}\oplus\sigma_{\ker\rho}\left(\log(\sigma_{\operatorname{supp}\rho}\oplus\sigma_{\ker\rho})-\log(\rho_{\operatorname{supp}\rho}\oplus\rho_{\ker\rho})\right)\}$
$=\operatorname{Tr}\{\sigma_{\operatorname{supp}\rho}(\log\sigma_{\operatorname{supp}\rho}-\log\rho_{\operatorname{supp}\rho})\}+\underbrace{\operatorname{Tr}\{\sigma_{\ker\rho}(\log\sigma_{\ker\rho}-\log\rho_{\ker\rho})\}}_{=0}$
$=S(\sigma_{\operatorname{supp}\rho}\|\rho_{\operatorname{supp}\rho}).$
Thus, we can replace $\mathcal{H}$ in Problem 2.4 by $\operatorname{supp}\rho$, and the operators by their restrictions to $\operatorname{supp}\rho$. Note that $\rho_{\operatorname{supp}\rho}$ is positive definite.
Step 2. Next we obtain the form of $\sigma_{\operatorname{supp}\rho}$. For ease of presentation let us simply denote $(\sigma/\rho/H_{i})_{\operatorname{supp}\rho}$ by $(\sigma/\rho/H_{i})$. With $\rho$ now positive definite, $\log\rho$ is well-defined. Now we invoke Proposition B.1 to extract the optimal $\sigma$ by setting $\frac{\partial\mathcal{L}}{\partial\sigma}=\mathbf{0}$.
Set up the Lagrangian
$\mathcal{L}=\operatorname{Tr}\{\sigma(\log\sigma-\log\rho)\}-\sum_{i}\lambda_{i}(\operatorname{Tr}(\sigma H_{i})-m_{i})-\eta(\operatorname{Tr}\sigma-1)$ (2.6)
where $\lambda_{i}$ and $\eta$ are the Lagrange multipliers. Making use of Propositions B.2 and B.3, setting $\frac{\partial\mathcal{L}}{\partial\sigma}$ to zero gives
$\implies\sigma=e^{\eta-1}e^{\lambda\cdot H+\log\rho}$
Step 3. Now we show that $\sigma^{\star}$ as given in Eq. 2.4 indeed minimizes $S(\sigma\|\rho)$. But this follows easily from Lemma 2.7. Furthermore, since $S(\sigma\|\rho)$ is a strictly convex functional of $\sigma$, it can have at most one minimizer in the convex set $M$, thereby showing the uniqueness of $\sigma^{\star}$. Finally, again by Lemma 2.7 we note that $\lambda^{\star}$ satisfies $\lambda^{\star}=\operatorname*{argmax}_{\lambda\in\mathbb{R}^{d}}\left[\lambda\cdot m-\log\operatorname{Tr}(e^{\lambda\cdot H+\log\rho})\right]=\operatorname*{argmin}_{\lambda\in\mathbb{R}^{d}}\log\operatorname{Tr}(e^{\lambda\cdot(H-m)+\log\rho})=\operatorname*{argmin}_{\lambda\in\mathbb{R}^{d}}\operatorname{Tr}(e^{\lambda\cdot(H-m)+\log\rho})$, where the last equality holds because $\log f(x)$ and $f(x)$ share the same minimum/maximum points, provided $f(x)>0$ at those points. ∎
Lemma 2.6.
$\sigma=\sigma_{\operatorname{supp}\rho}\oplus\sigma_{\ker\rho}.$
Proof.
[…]
The following lemma is the quantized version of Lemma A.1. We employ analogous arguments and notation, starting with
Lemma 2.7.
$S(\sigma\|\rho)\geq\sup_{\lambda\in\mathbb{R}^{d}}\left[\lambda\cdot m-\log\operatorname{Tr}(e^{\lambda\cdot H+\log\rho})\right].$ (2.7)
$S(\sigma_{\lambda^{\prime}}\|\rho)=\lambda^{\prime}\cdot m-\log\operatorname{Tr}(e^{\lambda^{\prime}\cdot H+\log\rho})=\sup_{\lambda\in\mathbb{R}^{d}}\left[\lambda\cdot m-\log\operatorname{Tr}(e^{\lambda\cdot H+\log\rho})\right].$ (2.8)
Proof.
$S(\sigma\|\rho)$ $=$ $S(\sigma\|\sigma_{\lambda})+\operatorname{Tr}\{\sigma(\log\sigma_{\lambda}-\log\rho)\}$ (2.9)
Motivated by the form of the state $\sigma_{\operatorname{supp}\rho}^{\star}$ in Theorem 2.5, we make the following definition:
Definition 2.8 (Quantum Esscher Transform).
@@
Remark 2.9.
The state $\sigma_{\operatorname{supp}\rho}^{\star}$ in Theorem 2.5 is thus a $(\lambda^{\star},H_{\operatorname{supp}\rho})$-quantum Esscher transform of $\rho_{\operatorname{supp}\rho}>0$. Also note that the quantum Esscher transform subsumes the classical Esscher transform as a special case, wherein $\rho,H_{i}$ are diagonal and thus commute.
2.2.3 Connection to quantum imaginary time evolution
Quantum imaginary-time evolution (QITE) is a conceptual tool which relates to the finding of ground states of Hamiltonians [${}^{+}$19, ${}^{+}$20]. From the real-time Schrödinger equation one obtains the imaginary-time Schrödinger equation $\frac{\partial|\psi\rangle}{\partial\tau}=-H|\psi\rangle$ by performing a Wick rotation, i.e. $\tau=it$. For general mixed states $\rho$, the imaginary-time Liouville-von Neumann equation [BK91] is given by
$\frac{\partial\rho}{\partial\tau}=-\{H,\rho\}+2\langle H\rangle\rho,$ (2.10)
@@ 2.2.3 Connection to quantum imaginary time evolution
In [OP07] it was asserted that under certain conditions, namely ‘when the prior and posterior states are close to each other with respect to the Fisher information metric’, the minimizing relative entropy problem could be solved by formally integrating a ‘quantum trajectory’ equation [OP07, Bra96]. This equation takes on the same form as Eq. 2.10, and thus its solution is given by Eq. 2.11. More specifically, we have
@@
where $\theta$ are the Lagrange multipliers. Here we simply observe that $\rho(\theta)$ resembles the imaginary-time-evolved state in Eq. (2.11) if $\theta$ is one-dimensional and after making the substitution $\tau=-\theta/2$. Since the quantum Esscher transform provides an exact solution to the problem, under the aforementioned condition we note the connection between the quantum Esscher transform and QITE.

### v2

@@ Abstract
I Introduction
In probability and statistics, it is often important to find low relative-entropy distributions from a given fixed distribution. In addition, further constraints, the form and interpretation of which depend on the problem at hand, are frequently imposed on the target distribution. An interesting example is the following: consider the process of inferring probability distributions from a set of measurement data. The available data play the role of the constraints—they put restrictions on what the true distribution could be—but these data may not suffice to uniquely determine a probability distribution. In this situation, a common approach is to invoke Jaynes’ maximum entropy principle (MaxEnt) [18]. In essence, MaxEnt advocates that the selected distribution be the one that simultaneously maximizes entropy and satisfies the given constraints.
However, the situation becomes more nuanced if we already possess some knowledge of the system, say a prior distribution. In such cases, there is a more refined strategy: the minimum relative entropy principle. As expounded in [35, 32, 38], this principle, regarded as a generalization of MaxEnt, operates by minimizing the distinguishability (characterized by the relative entropy) between the prior distribution and the distribution to be selected, while respecting the imposed constraints. This systematic approach to incorporating new data makes it fundamental in Bayesian statistics. The updating procedure results in the posterior distribution which reflects the most current understanding of the system in light of the observed data.
When the measurement data is presented in the form of expectation values of selected random variables, the solution to the corresponding relative entropy minimization problem takes the form known as an Esscher Transform. Named after Swedish mathematician and economist Fredrik Esscher, who introduced the concept in 1932 in his work on risk theory [8], the Esscher Transform, also known as ‘exponential tilting’ in statistics, and its various extensions have since then found many applications beyond minimizing relative entropy. Notable examples include option pricing (in mathematical finance) [14], importance sampling (for rare-event simulation) [34] and Lévy processes (in financial economics) [17]. More recently, it has also made inroads into machine learning [5], in the context of empirical risk minimization.
@@ I Introduction
$\text{minimize}_{\sigma\geq 0}$ $\;\;S(\sigma\|\rho)$ (I.1)
@@
where $\rho$ is the a priori state and $H_{i}$, $i\in[d]$ are observables. Refer to Definition 1 for the precise formulation. In the first part of this work, we show the formal solution to this constrained optimization problem. The solution methodology is modelled after its classical predecessor, albeit with added technical intricacies to manage. The form of the corresponding solution then motivates us to define the quantum Esscher Transform, see Definition 2. The proof of the solution to the optimization problem is found in Theorem 2. The quantum Esscher Transform can be viewed as a generalization of the (classical) Esscher Transform, and indeed subsumes the latter as a special case. In the second part of this work, with an eye toward potential applications, we discuss the implementation of the quantum Esscher Transform on fault-tolerant quantum computers. Our algorithm is based on the modern techniques of block-encoding and quantum singular value transformation (QSVT) [15, 27]. As an input model we consider purifications of the density operator $\rho$ and block-encodings of the operators $H_{i}$. The main algorithm is Algorithm 1, whose complexity is discussed in Theorem 5. The quantum Esscher Transform could find applications in quantum analogues of problems in statistics, machine learning, and finance.
I.1 Preliminaries and notation
@@
Figure 1: Overview of the paper: Given an input density operator $\rho$, constraint observables $H_{i}$ and parameters $\theta_{i}$, $i\in[d]$, the quantum Esscher Transform of $\rho$ is given by $\sigma:=e^{\theta\cdot H+\log\rho}/Z$, where $Z=\operatorname{Tr}(e^{\theta\cdot H+\log\rho})$ is the normalization factor. Our algorithm QEsscher is based on block-encodings and QSVT. It takes as inputs block-encodings of $\rho$, $H_{i}$, and the parameter vector $\vec{\theta}$ and outputs the block-encoding of $\sigma$. The state $\sigma$ itself can also be obtained using Gibbs state preparation techniques in a subsequent step. Finally, we envision potential applications of the quantum Esscher Transform in quantum analogues of problems in statistics, machine learning, and finance.
II Quantum Esscher Transform
II.1 Esscher Transform
@@
The Esscher Transform is a map from and onto the space of probability mass/density functions, as $\mathcal{E}(f;\theta)=f_{\theta}$. In this work, we never invoke $\mathcal{E}$ and simply call $f_{\theta}$ the Esscher Transform of $f$, in the same spirit as the Fourier Transform. In the context of probability theory, let $(\Omega,\Sigma,P)$ be a probability space and $X:\Omega\longrightarrow\mathbb{R}^{d}$ a random $d$ dimensional vector. This setting motivates the equivalent definition (see Remark 2 below) of Esscher Transforms for measures/distributions.
Definition 1 (Esscher Transform for probability distributions).
@@
$\text{minimize}_{Q\in[0,1]^{|\Omega|}}\;\;$ $D(Q\|P)$ (II.1)
@@
Note that there are $d+1$ constraints on $Q$, hence in feasible, non-redundant cases we have $d+1\leq|\Omega|$, or equivalently $d<|\Omega|$. We have the following solution to the optimization problem.
Theorem 1.
Given a random vector $X:\Omega\longrightarrow\mathbb{R}^{d}$ and $m\in\mathbb{R}^{d}$ where $\min_{\omega\in\Omega}X_{i}(\omega)<m_{i}<\max_{\omega\in\Omega}X_{i}(\omega)$ for $i\in[d]$ where $d<|\Omega|$. There exists a unique solution $Q^{\star}$ to problem II.1, given by
@@
where $\lambda^{\star}:=\operatorname*{argmin}_{\lambda\in\mathbb{R}^{d}}\mathbb{E}_{P}[e^{\lambda\cdot(X-m)}]$. Thus $Q^{\star}$ is the $(\lambda^{\star},X)$-Esscher Transform of $P$, see Definition 1.
The proof is elaborated in Appendix A. Let us comment on a subtlety. Above, we have called $Q^{\star}$ the Esscher Transform of $P$. Recall that the Esscher Transform as originally defined by Esscher pertains to probability mass/density functions instead of measures. In Remark 2, we show that using the same terminology for probability measures is well-justified, at least for the case when $\Omega$ is discrete.
II.2 Quantum version
Problem statement -- Many concepts in classical probability theory have meaningful generalizations in quantum theory. For example, sample spaces, probability distributions and random variables find their respective counterparts in Hilbert spaces, density operators and observables (the latter also include the former as special instances). The quantum counterpart of the relative entropy is the quantum relative entropy,
@@ II.2 Quantum version
defined for density operators $\sigma,\rho$. As in the classical case, the definition of $S(\sigma\|\rho)$ imposes constraints on $\sigma$ and $\rho$ in order to have $S(\sigma\|\rho)<\infty$. Namely, $\operatorname{supp}(\sigma)\subseteq\operatorname{supp}(\rho)$ (see Chapter 11, [37]) or equivalently, $\ker(\rho)\subseteq\ker(\sigma)$. Using terminology from measure theory, if this condition is satisfied we say $\sigma$ is absolutely continuous with respect to $\rho$ ($\sigma\ll\rho$). This is analogous to the absolute continuity between probability distributions in classical probability theory. Now we formally state the quantized version of Problem II.1.
Problem 1.
Let $\mathcal{H}_{N}$ be an $N$-dimensional Hilbert space and $\rho\in\mathcal{D}(\mathcal{H}_{N})$ be a density operator. For $i\in[d]$ where $d<N^{2}$, let $H_{i}$ be an observable with $h_{i,\min}$ and $h_{i,\max}$ denoting its smallest and largest eigenvalue respectively. For $m\in\mathbb{R}^{d}$ with $h_{i,\min}<m_{i}<h_{i,\max}$, solve
$\text{minimize}_{\sigma\geq 0}\;\;$ $S(\sigma\|\rho)$ (II.2)
@@
Solution -- Before considering the solution, let us briefly comment on a few possible concerns. First, $S(\sigma\|\rho)$ requires taking the logarithm of $\rho$, which poses a problem if $\rho$ is not strictly positive definite. This issue is circumvented if, as mentioned above, $\ker(\rho)\subseteq\ker(\sigma)$. The analysis becomes relatively straightforward if we partition the Hilbert space $\mathcal{H}$ into suitable subspaces and examine $\sigma$ over them separately. To this end, we introduce the following notation. Let $\mathcal{G}$ be a subspace of $\mathcal{H}$. For $A\in\mathcal{L}(\mathcal{H})$, denote $A_{\mathcal{G}}:=\Pi_{\mathcal{G}}A\Pi_{\mathcal{G}}\in\mathcal{L}(\mathcal{G})$, where $\Pi_{\mathcal{G}}$ is the projector onto $\mathcal{G}$.
@@
We have the following result, which partially resolves Problem 1:
Theorem 2.
The solution to Problem 1 takes the form
$\sigma^{\star}=\sigma_{\operatorname{supp}\rho}^{\star}\oplus\sigma_{\ker\rho}^{\star},$ (II.3)
@@
$\sigma_{\operatorname{supp}\rho}^{\star}=\frac{e^{\lambda^{\star}\cdot H_{\operatorname{supp}\rho}+\log\rho_{\operatorname{supp}\rho}}}{\operatorname{Tr}(e^{\lambda^{\star}\cdot H_{\operatorname{supp}\rho}+\log\rho_{\operatorname{supp}\rho}})}\qquad\text{and}\qquad\sigma_{\ker\rho}^{\star}=\mathbf{0}.$ (II.4)
@@
$\operatorname{Tr}\left(e^{\lambda^{\star}\cdot(H_{\operatorname{supp}\rho}-m)+\log\rho_{\operatorname{supp}\rho}}(H_{i,\operatorname{supp}\rho}-m_{i})\right)=0,i\in[d].$ (II.5)
The proof of Theorem 2 draws inspiration from the classical version but additional technical hurdles need to be overcome, as mentioned above. The details are elaborated in Appendix C.
Motivated by the form of the state $\sigma_{\operatorname{supp}\rho}^{\star}$ in Theorem 2, we make the following definition:
Definition 2 (Quantum Esscher Transform).
@@
Remark 1.
The state $\sigma_{\operatorname{supp}\rho}^{\star}$ in Theorem 2 is thus a $(\lambda^{\star},H_{\operatorname{supp}\rho})$-quantum Esscher Transform of $\rho_{\operatorname{supp}\rho}>0$. Also note that the quantum Esscher Transform subsumes the classical Esscher Transform as a special case, wherein $\rho,H_{i}$ are diagonal and thus commute.
Connection to quantum imaginary time evolution -- Quantum imaginary-time evolution (QITE) is a conceptual tool used to find ground states of Hamiltonians [26, 28]. From the real-time Schrödinger equation one obtains the imaginary-time version $\frac{\partial|\psi\rangle}{\partial\tau}=-H|\psi\rangle$ by performing a Wick rotation, i.e. setting $\tau=it$. For general mixed states $\rho$, the imaginary-time Liouville-von Neumann equation [3] is given by
$\frac{\partial\rho}{\partial\tau}=-\{H,\rho\}+2\langle H\rangle\rho,$ (II.6)
@@
In [32] it was asserted that under certain conditions, namely ‘when the prior and posterior states are close to each other with respect to the Fisher information metric’, the minimizing relative entropy problem could be solved by formally integrating a ‘quantum trajectory’ equation [32, 4]. This equation takes on the same form as Eq. II.6, and thus its solution is given by Eq. II.7. More specifically, we have
@@
where $\theta$ are the Lagrange multipliers. Here we simply observe that $\rho(\theta)$ resembles the imaginary-time-evolved state in Eq. (II.7) if $\theta$ is one-dimensional and after making the substitution $\tau=-\theta/2$. Since the quantum Esscher Transform provides an exact solution to Problem 1, under the aforementioned conditions presumed by [32], we note the connection between the quantum Esscher Transform and QITE.
Figure 2: Relative entropy between the quantum Esscher transformed state $\rho_{\text{Esscher}}$ and the imaginary-time evolved $\rho_{\text{QITE}}$ as a function of the parameter $\theta$. Consider a single qubit system, with initial positive state $\rho=0.5(|0\rangle\langle 0|+|1\rangle\langle 1|)+a(|0\rangle\langle 1|+|1\rangle\langle 0|)$, where we vary $a$ from $0.0$ to $0.4$. We also choose the Hamiltonian $H=\sigma_{x}+\sigma_{z}$. The plot above illustrates how the distinguishability of these two states varies with respect to $\theta$.
In Fig. 2 we provide a simple plot illustrating the difference between $\rho_{\text{Esscher}}$ and the $\rho_{\text{QITE}}$ for a simple initial state $\rho=0.5(|0\rangle\langle 0|+|1\rangle\langle 1|)+a(|0\rangle\langle 1|+|1\rangle\langle 0|)$, where $a\in[0,0.5)$ is the coherence parameter, and Hamiltonian $H=\sigma_{x}+\sigma_{z}$. We use relative entropy as the measure of distinguishability between the two states. We note that as $\theta\rightarrow 0$, the relative entropy vanishes, as is immediate from the definitions. We also note that as $\theta\rightarrow\pm\infty$ the relative entropy vanishes – because the $\theta\cdot H$ terms then dominate the $\log\rho$ term. In intermediate ranges for $\theta$, the relative entropy is nonzero, and this is interpreted as the error arising when one uses $\rho_{\text{QITE}}$ as a proxy for $\rho_{\text{Esscher}}$. For the dependency on $a$, since $a$ parametrizes the noncommutativity between $\rho$ and $H$, when $a\to 0$ the relative entropy remains small as $\theta$ varies. The relative entropy becomes maximal when $a\rightarrow 0.5$.